OpenStudy (anonymous):

never fully understood these problems. Please help! y varies directly as the square of x and inversely as the square root of z. what is the effect on y if x is halved and z is quadrupled?

OpenStudy (anonymous):

i know there is supposed to be a k somewhere in the equation but so far i have \[yx ^{2}/\sqrt{z}\]

OpenStudy (asnaseer):

ok, best to look at this in parts. start with the statement "y varies directly as the square of x" do you know what that means?

OpenStudy (anonymous):

it means multiply right?

OpenStudy (asnaseer):

it "means" y changes as some constant times "x squared" we would normally write this as:\[y=kx^2\]where k is some constant.

OpenStudy (asnaseer):

understand so far?

OpenStudy (anonymous):

ohh ok. got it so far.

OpenStudy (anonymous):

so would it be? y=\[kx ^{2}/\sqrt{z}\]

OpenStudy (asnaseer):

good, now lets look at the second part: "y varies inversely as the square root of z" this would imply that y changes as some constant divided by the square root of z, and we would write this as:\[y=\frac{k_2}{\sqrt{z}}\]where \(k_2\) is just some other constant.

OpenStudy (asnaseer):

yes - you have it right, if we put these two statements together, then we get:\[y=\frac{kx^2}{\sqrt{z}}\]where k is some constant

OpenStudy (asnaseer):

now what would happen to y if x is halved?

OpenStudy (anonymous):

we would take the square root since its \[x ^{2}\] right?

OpenStudy (asnaseer):

no, you need to replace the 'x' in the equation with 'x/2' and see what you get

OpenStudy (anonymous):

sorry this is the part i get confused…ok so we do \[(x/2)/\sqrt{4}\]?

OpenStudy (anonymous):

but I'm not understanding what we do with the y

OpenStudy (asnaseer):

ok, let me try to explain it to you step-by-step...

OpenStudy (anonymous):

ok thank you.'

OpenStudy (asnaseer):

we have:\[y=\frac{kx^2}{\sqrt{z}}\]now lets replace x with x/2 to get a new y as follows:\[y_{new}=\frac{k(x/2)^2}{\sqrt{z}}=\frac{kx^2/4}{\sqrt{z}}=\frac{kx^2}{4\sqrt{z}}\]now notice that this is a quarter of the original y. So halving x has reduced y by a quarter.

OpenStudy (asnaseer):

does that make sense?

OpenStudy (anonymous):

its still a little bit confusing to me but since you showed me step by step i understand it better now. Thank you for your help!

OpenStudy (asnaseer):

no problems my friend. now lets see what happens to y if z is also quadrupled...

OpenStudy (asnaseer):

we know that if we halve x we end up with:\[y_{new}=\frac{kx^2}{4\sqrt{z}}\]now lets replace z with 4z and see what happens to y now:\[y_{new}=\frac{kx^2}{4\sqrt{4z}}=\frac{kx^2}{4\times2\sqrt{z}}=\frac{kx^2}{8\sqrt{z}}\]now compare this to the original y which was given by:\[y=\frac{kx^2}{\sqrt{z}}\]how has y changed when we halved x and quadrupled z?

OpenStudy (anonymous):

it becomes ⅛ right? since before z was quadrupled it was ¼ and the square root of 4 is 2 so we multiply 2 and 4

OpenStudy (asnaseer):

perfect! I think you are starting to understand this better now. :)

OpenStudy (anonymous):

good! i feel like I am too! thank you so much for your help!

OpenStudy (asnaseer):

yw :)